OpenStudy (moongazer):

12 years from now Maria will be twice as old as Steven. The current sum of the ages of Maria and Steven is 36. How old is Steven right now?

OpenStudy (moongazer):

Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (anonymous):

let maria be of x years steven be of y years x+12=2y (x+12)/2=y x+y=36 y=36-x y=(x+12)/2 36-x=(x+12)/2 2(36-x)=x+12 72-2x= x+12 60=3x 60/3=x 20=x x+y=36 20+y=36 y=16 current ages ; maria :20 steven:16

OpenStudy (anonymous):

hope i am right moongazer

OpenStudy (moongazer):

look at this

OpenStudy (moongazer):

Problem Correct Answer Your Answer 1 12 years from now Maria will be twice as old as Steven. The current sum of the ages of Maria and Steven is 36. How old is Steven right now? Solution Maria + 12 = 2 × (Steven + 12) Maria + Steven = 36 Maria = 36 - Steven 36 - Steven + 12 = 2 × (Steven + 12) 48 - Steven = 2 × Steven + 24 48 = 3 × Steven + 24 3 × Steven = 48 - 24 = 24 Steven = 8

OpenStudy (moongazer):

i got that from a website

OpenStudy (moongazer):

i can't understand it

OpenStudy (anonymous):

oh sorry did mistake sorry

OpenStudy (moongazer):

explain please

OpenStudy (moongazer):

why it is the answer

OpenStudy (anonymous):

what u want not answer is already explained

OpenStudy (moongazer):

i am just confused

OpenStudy (anonymous):

x+12=2(y+12) x+y=36 x+12=2y+24 x+12-24=2y x-12=2y (x-12)/2=y x+y=36 x+(x-12)/2=36 (2x+x-12)/2=36 3x-12=72 3x=84 x=28 y=(28-12)/2=16/2=8 years

OpenStudy (anonymous):

got it now moongazer

OpenStudy (moongazer):

what does this mean? 12 years from now Maria will be twice as old as Steven

OpenStudy (moongazer):

does it mean that if you will multiply the age of Steven to 2 after 12 years the age of Maria will be equal to the age of steven?

OpenStudy (moongazer):

i am not that good in english :)

OpenStudy (anonymous):

this means let maria's age be x and steven's age be y it says that 12 years from now x+12,y+12 x+12=twice of y+12 x+12=2(y+12)

OpenStudy (anonymous):

got it now moongazer well if u are on facebook then tell me let me explain u there more

OpenStudy (moongazer):

so if i say, 5 years from now mark's age will be twice as mario's age does this mean after 5 years mark's age will be equal to mario's if i will multiply marios age by 2?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

u on facebook moongazer

OpenStudy (moongazer):

give me other example of that kind of statement

OpenStudy (moongazer):

yup

OpenStudy (moongazer):

i am playing tetris

OpenStudy (anonymous):

give me ur profile link

OpenStudy (moongazer):

please give me another example

OpenStudy (anonymous):

after 12 years i would be twice as ur's age let my age be x ur's age be y x+12=2(y+12)

OpenStudy (moongazer):

i get it know thanks!

OpenStudy (anonymous):

my pleasure give me ur profile link of facebook

OpenStudy (moongazer):

why?

OpenStudy (moongazer):

give me your first i don't trust anyone at the internet sorry

OpenStudy (anonymous):

facebook.com/kushashwaravi.shrimali

OpenStudy (anonymous):

m = 20, s = 16